# A Reliable Framework for Human-in-the-Loop Anomaly Detection in Time Series

### v2

@@ III-D Performance Evaluation Metrics
To assess the participants’ subjective perception of human–machine cooperation, a questionnaire with a five-point Likert Scale [24] with the following items was used.
Figure 4: Tukey HSD post-hoc test results for MSL dataset. The left plot illustrates classification performance, while the right one depicts attention accuracy. Methods not sharing any letter are significantly different by the Tukey-test at the 5% level of significance.
III-E Datasets
Mars Science Laboratory (MSL) dataset is a dataset that corresponds to the sensor and actuator data for the Mars rover [26]. At the same time, the data set marks the time point when the abnormal behavior occurs for us to calculate relevance accuracy. However, this dataset is known to have many trivial sequences [32]; hence, to make a best-faith effort to filter out data with mislabeled ground truth time points, we build the new dataset based on the two nontrivial ones (A2 and A4), where we randomly divide the original time series into 1000 equal-length sequences, all of 800 lengths, where the ratio of labeled normal to abnormal data is 3 to 2.
The Simulation Testbed for Exploration Vehicle ECLSS (STEVE) [33] at the University of Colorado Boulder is a simplified single-bed $CO_{2}$ removal system of the Carbon Dioxide Removal Assembly (CDRA) onboard the International Space Station (ISS). A corresponding Simulink model is provided to simulate multiple failure modes similar to the STEVE testbed. In our study, we conducted a simulation consisting of four cycles, each comprising 80 minutes of $CO_{2}$ adsorption followed by 80 minutes of desorption. During the third cycle, a simulated leak failure is introduced at a random time point in Valve 1, resulting in observable anomalies in three sensors: Bed Temperature, $CO_{2}$ Concentration, and Flow Rate. These sensor sequences are treated collectively as a set of sequences. We generated 1000 sets of sequences with a ratio of 4 to 1 between labeled normal and abnormal data. Additionally, we marked the time points at which abnormal behaviors occurred in the three sensors to calculate relevance accuracy.
III-F Baseline Methods
Baseline model. In this paper, we use the fully convolutional network (FCN) proposed in [34], which has demonstrated good performance on time series classification benchmarks. In addition to the baseline model, we also compare our spurious attention elimination method with the following data augmentation strategies for the same number, $Q$, of instances:
Random Aug. Randomly select $Q$ instances from the training set for model fine-tuning.
Random Mask. Select the $Q$ instances proposed by our system, but add masks at random time points for each instance. To further verify the effectiveness of our method, we compared the following strategy as the ablation study of our method.
Cluster Mask. The $Q$ instances proposed by our system only concede the participant’s operation in the Cluster Information window, which means the participants in the Instance Information window make no further adjustments. This method uses the same masking strategy as ours.
III-G Statistical Analysis
For Hypothesis 1, since all five methods involved in the comparison have been tested an equal number of times, we directly assess their performance using the metrics introduced in Section III-D. Therefore, our hypothesis tests reduces to:
@@ III-G Statistical Analysis
IV Results
@@
Figure 5: Tukey HSD post-hoc test results for STEVE dataset. The left plot illustrates classification performance, while the right one depicts attention accuracy. Methods not sharing any letter are significantly different by the Tukey-test at the 5% level of significance.
Figure 6: Participants perception of the system, according to nine Likert-Type Questions using a 5-point scale.
@@
In order to effectively measure the objective performance difference between purely automatic algorithm-driven and human-machine cooperative anomaly detection methods, we conducted an analysis using Analysis of Variance (ANOVA) tests [35]. Objective performance is divided into two aspects, namely classification correctness and attention correctness. For the former, we utilize the ANOVA test to compare the means of accuracy, precision, recall, and F1 score across different models to statistically ascertain any significant performance variations. For the latter, we employed the ANOVA test to determine whether there are statistically significant differences in relevance accuracy at varying levels of $M$, specifically at $M=20,25,30$, and $35$. Upon identifying significant differences with the ANOVA test, we proceeded with the Tukey HSD post-hoc test to further distinguish which specific group means, among the compared groups, are significantly different from each other, ensuring a thorough evaluation of pairwise contrasts. Owing to space limitations, this article presents only the results of the Tukey HSD post-hoc test for two datasets, as illustrated in Fig. 4 and Fig. 5, respectively.
@@
The result of the five-point Likert scale for objective evaluation is shown in Fig. 6. For a more comprehensive analysis, we also furnish the post-test results for the three methods, assessing their comparative effectiveness using a t-test. The stars intended to flag levels of significance as shown in Fig. 6 are based on the adjusted p-values.
V Discussion
V-A Hypothesis 1
@@
V-B Hypothesis 2
In the subjective evaluation, both the cluster mask method and our method are significantly different from the automatic algorithm in intuition, transparency, and reliability. Because the automatic algorithm can only provide the result of classification, can not provide the criterion of classification can not interact. The cluster mask method is not significantly different from intuition and transparency, as both provide reasons for the model to make a prediction. But on reliability, we are significantly different from the cluster mask method because of the error of the automatic clustering algorithm itself, while in our method, the D window can provide more detailed options to further correct the automatic algorithm, which helps improve performance objectively. At the same time, humans are subjectively more likely to believe that our methods are more reliable.

### v4

@@ 4.6 Performance Evaluation Metrics
To assess the participants’ subjective perception of human–machine cooperation, a questionnaire with a five-point Likert Scale De Winter and Dodou (2010) with the following items was used.
4.7 Statistical Analysis
For Hypothesis 1, since all five methods involved in the comparison have been tested an equal number of times, we directly assess their performance using the metrics introduced in Section 4.6. Therefore, our hypothesis tests reduces to:
@@ 4.7 Statistical Analysis
4.8 Annotation Efficiency Evaluation Settings
We quantify and compare the human effort between HILAD and multiple representative baselines that also utilize human annotations. Specifically, our comparative study involves three active learning baselines: random sampling, uncertainty-based sampling, and diversity-based sampling. To align with prior practices, we view our basic model as the warm-up model that is trained with labeled data and might have spurious issues. Then, during the human interaction phase, we present the user with individual time sequences with CAM highlighting, and ask them to select from binary options, “correct” or “spurious”, indicating whether spurious correlation exists. View D in the HILAD interface is used as the annotation interface. We fix the annotation budget as 160, which is 20% of the total instances. Additional settings are as follows.
  • Random-based sampling. Instances are randomly sampled from the dataset and listed in the interface for user annotation. This reflects real-world, unstructured manual debugging workflows.
  • Uncertainty-based sampling. Individual instances are ranked in descendingly order by model uncertainty (e.g., based on softmax entropy). Users are given instances in order and assign spuriousness labels to each of them.
  • Diversity-based sampling. We use the basic setting to cluster data features and sample instances nearest to each cluster centroid. Then present the instances to the user in a random order.
Considering the practical challenges, we re-invite five of our participants in the original study. We asked them to perform annotations following four settings: ours, random, uncertainty, and diversity. Their orders are random to reduce familiarity-caused noise, and we have a 10-minute break in between. Note that for HILAD, the users are asked to determine how many instances to annotate, according to their interactions and observations through our interface.
5 Results
@@
Figure 4: Tukey HSD post-hoc test results for MSL dataset using FCN. The left and right plots illustrate classification and attention accuracy, respectively. Methods not sharing any letter are significantly different by the Tukey-test at the 5% level of significance. Note: y-axes are truncated to emphasize fine-grained differences.
Figure 5: Tukey HSD post-hoc test results for STEVE dataset using FCN. The left and right plots illustrate classification and attention accuracy, respectively. Methods not sharing any letter are significantly different by the Tukey-test at the 5% level of significance. Note: y-axes are truncated to emphasize fine-grained differences.
5.1 Quantitative evaluation
@@
In order to effectively measure the objective performance difference between purely automatic algorithm-driven and human-machine cooperative anomaly detection methods, we conducted an analysis using Analysis of Variance (ANOVA) tests St et al. (1989). Objective performance is divided into two aspects, namely classification correctness and attention correctness. For the former, we utilize the ANOVA test to compare the means of accuracy, precision, recall, and F1 score across different models to statistically ascertain any significant performance variations. For the latter, we employed the ANOVA test to determine whether there are statistically significant differences in relevance accuracy at varying levels of $M$, specifically at $M=20,25,30$, and $35$. Upon identifying significant differences with the ANOVA test, we proceeded with the Tukey HSD post-hoc test to further distinguish which specific group means, among the compared groups, are significantly different from each other, ensuring a thorough evaluation of pairwise contrasts.
We report the Tukey HSD post-hoc test results on the MSL and STEVE datasets using FCN in Figs 4 and 5. For the more challenging real-world KPI dataset, we include both FCN and TST models, with results shown in Figs. 6 and 7. Across all settings, HILAD consistently outperforms baseline and augmentation-based methods in both prediction accuracy and attribution alignment. These results underscore the importance of evaluating not only classification outcomes but also the underlying rationale, reinforcing the need for human-in-the-loop validation.
5.2 Annotation efficiency evaluation
The results are reported in Table 1, which includes the average annotation time, annotated data size, the spuriousness label coverage (how many instances have been assigned a label), and the F1 gain in percentage compared to the basic model. We can observe that HILAD consistently achieved the best trade-off: requiring the least annotation time (8.3 $\pm$ 2.1 min), while delivering the highest F1 improvement (18.485 $\pm$ 2.7%) and covering over 800 sequences thanks to our propagation design. In contrast, despite much larger annotation budgets being used, the active learning baselines still underperformed because they solely rely on manual annotations, leading to significantly less spuriousness label coverage. These results demonstrate that our framework supports minimal but essential human input. With fewer human annotations, our design of attribution-guided behavior summarization and annotation propagation leverages human efforts at a lower cost, which is not achievable through unstructured or conventional data sampling.
Figure 6: Tukey HSD post-hoc test results for KPI dataset using FCN. The left and right plots illustrate classification and attention accuracy, respectively. Methods not sharing any letter are significantly different by the Tukey-test at the 5% level of significance. Note: y-axes are truncated to emphasize fine-grained differences.
Figure 7: Tukey HSD post-hoc test results for KPI dataset using TST. The left and right plots illustrate classification and attention accuracy, respectively. Methods not sharing any letter are significantly different by the Tukey-test at the 5% level of significance. Note: y-axes are truncated to emphasize fine-grained differences.
Method Avg. annotation time (minutes) Avg. annotation number Coverage F1 Gain (%)
random 14.2 $\pm$ 1.8 160 160 1.908 $\pm$ 1.2
uncertainty 13.3 $\pm$ 1.4 160 160 2.610 $\pm$ 0.4
diversity 14.7 $\pm$ 2.3 160 160 2.328 $\pm$ 1.9
HILAD (ours) 8.3 $\pm$ 2.1 10.2 $\pm$ 3.4 800 18.485 $\pm$ 2.7
Table 1: Comparison of annotation time, number, coverage, and resulting F1 gain across HILAD and baseline annotation strategies on MSL dataset.
5.3 Objective evaluation
@@
The result of the five-point Likert scale for objective evaluation is shown in Fig. 8. For a more comprehensive analysis, we also furnish the post-test results for the three methods, assessing their comparative effectiveness using a t-test. The stars intended to flag levels of significance, which are based on the adjusted p-values.
5.4 Ablation study
To validate our design choice of $\alpha$ = 0.5 in the joint similarity computation for behavior clustering, we conducted a sensitivity analysis by varying $\alpha$ across a range of values. Refer to Formula 4 in Sec. 3.2.2, $\alpha$ controls the balance between time sequence similarity (via DTW) and attribution similarity (via CAM), with higher $\alpha$ favoring time sequence similarity. We first vary $\alpha$ across $\{0.0,0.5,1.0\}$ and provide qualitative observations on the produced clusters in Fig. 9. Specifically, in each setting, we randomly sample three clusters from the results and observe their included instances. The results show $\alpha=0.0$ overlooks the data consistency and $\alpha=1.0$ contrarily misses the model attribution consistency, while our choice $\alpha=0.5$ produces clusters with internally consistent time sequences and model attributes simultaneously.
Figure 8: Participants perception of the system, according to Likert-Type Questions using a 5-point scale.
We also quantitatively evaluate the impact of $\alpha$ on model performance, where we utilize FCN as the baseline on MSL data, measuring classification performance and attention correctness. In the experiment, we only allow cluster-level annotation and disable the adjustment of individual instances to ensure valid evaluation of $\alpha$. As shown in Table 2, $\alpha=0.5$ consistently produces the highest F1 and attention scores, supporting the use of equal weighting between sequence context and model behavior. Extreme values (e.g., $\alpha=0.0$ or $1.0$) led to performance drops, where $\alpha=1.0$ (only consider data similarity) results in the worst performance, as we lost neighbor consistency regarding model attributions. This experiment validates our parameter choice, confirming that collectively considering both temporal patterns and model attention is critical for meaningful behavior grouping.
6 Discussion
6.1 Hypothesis 1
@@
Figure 9: Qualitative observation on randomly sampled clusters when using different $\alpha$ values. We can observe that $\alpha=0$ leads to clusters with more coherent model attention only, while $\alpha=1$ produces clusters with consistent time sequences but dissimilar model behavior patterns. In contrast, our choice of $\alpha=0.5$ provides clusters with both consistent model attribution and similar time sequences.
$\alpha$=0.00 $\alpha$=0.25 $\alpha$=0.50 $\alpha$=0.75 $\alpha$=1.0
F1 88.21 $\pm$ 0.26 89.92 $\pm$ 0.27 90.62 $\pm$ 0.25 88.33 $\pm$ 0.60 87.1 $\pm$ 0.76
ATT 86.43 $\pm$ 0.79 89.13 $\pm$ 0.78 89.90 $\pm$ 0.77 85.82 $\pm$ 0.85 85.66 $\pm$ 0.84
Table 2: Quantitative evaluation on the effect of the $\alpha$ parameter that controls the balance between temporal and attribution similarity in clustering, showing $\alpha=0.5$ achieves the best trade-off in both classification (F1) and attention (ATT) performance.
@@
6.2 Hypothesis 2
In the subjective evaluation, both the cluster mask method and our method are significantly different from the automatic algorithm in intuition, transparency, and reliability. Because the automatic algorithm only provides classification results, it neither reveals the underlying decision criteria nor supports interaction. The cluster mask method is not significantly different from intuition and transparency, as both provide reasons for the model to make a prediction. But on reliability, we are significantly different from the cluster mask method because of the error of the automatic clustering algorithm itself, while in our method, the D window can provide more detailed options to further correct the automatic algorithm, which helps improve performance objectively. At the same time, humans are subjectively more likely to believe that our methods are more reliable.